the weight of all male senior high school students from one province is measured and is found to be approximately normally distributed with a mean of 115 lbs and standard deviation of 3 lbs. find the percentile rank of the following weights?



A. 112 lbs



B. 115 lbs

1
Expert's answer
2022-03-08T07:11:02-0500

Population Mean "(\u03bc) =115"

Population Standard Deviation "(\u03c3) =3"


"Z=\dfrac{X-μ}{σ}"


(a)"Z=\dfrac{112-115}{3}"

"Z=-1.0"

From the standard normal distribution tables

"Z(-1.0)=0.15866"

=15.866th percentile


(b) "Z=\dfrac{115-115}{3}"

"Z=0"

From the standard normal distribution tables

"Z(0)=0.50000"

=50th percentile